x 3 +1=2 ( 3 √2x-1) The quantity in brackets is under cube root.

x3+1=2 (3√2x-1)


 


The quantity in brackets is under cube root.

Grade:upto college level

2 Answers

Sher Mohammad IIT Delhi
10 years ago

-2330_untitled.jpg
the only easy way to solve this problem is to try plot the curve of both the plots, and see where they intersects, from the curve it can be seen the solution are , 0, .6, and 1.


MuraliKrishna Medavaram
10 years ago
f(x)=(x^3+1)/2 and g(x)=(v2x-1)^(1/3) are inverse functions of each other f(x)=g(x) has a solution when f(x)=X BECAUSE IF THEY INTERSECT they intersect on y=x so we can have the roots for (x^3+1)/2=x as x= 1 and (-1+-root5)/2 Thanks and Regards, M.MURALIKRISHNA askIITians faculty